Find the complete solution of xdy + (x ^ 3 - xy ^ 2 - y) dy = 0

1
Expert's answer
2022-03-07T16:20:01-0500
"xdy+(x^3-xy^2-y)dy=0\\
(x^3-xy^2-y)dy=-xdy\\
(1-1/2-1/3)d[x^3y+xy^3+y^2]=-d[xy]\\
\intop{1/6d[x^3y+xy^3+y^2]}=\intop{-d[xy]}\\
\intop{d[x^3y+xy^3+y^2]}=\intop{-6d[]}\\
x^3y+xy^3+y^2=-6xy+C, C-const\\
x^3y+xy^3+y^2+6xy=C, C-const\\"

Answer: "x^3y+xy^3+y^2+6xy=C, C-const" and (x=0,y=0) is also a solution.